OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the graph at the given point. y^2 +ln(xy) = 2, (e, 1)

OpenStudy (anonymous):

start with \[2yy'+\frac{1}{xy}\left(y+xy'\right)=0\]

OpenStudy (anonymous):

Makes sense so far.

OpenStudy (anonymous):

first part is clear i hope, second is chain rule plus product rule now you have a choice: you can solve for \(y'\) using algebra, then replace \(x\) by \(e\) and \(y\) by \(1\) to get the slope, or you can make the replacement at this step and then solve for \(y'\)

OpenStudy (anonymous):

Could you show me the first way?

OpenStudy (anonymous):

i bunch of algebra, easier to do it with pencil and paper

OpenStudy (anonymous):

Alright, that is fine. I should be able to get it from here. Thank for the help!

OpenStudy (anonymous):

i can try distribute first \[2yy'+\frac{1}{x}+\frac{y'}{y}=0\] then \[2yy'+\frac{y'}{y}=-\frac{1}{x}\] \[y'(2y+\frac{1}{y})=-\frac{1}{x}\] then divide hope my algebra is good

OpenStudy (anonymous):

So then the tangent line would be (1-y) = (-1/x)/(2y+(1/y))(e-x) ?

OpenStudy (anonymous):

there should be no \(x\) or (y\) in your answer \(x=e,y=1\)

OpenStudy (anonymous):

oh is see, the equation for the line, sorry

OpenStudy (anonymous):

\[y-1=m(x-e)\] for whatever ugly expression you got for \(m\)

OpenStudy (anonymous):

but \(m\) is the slope, it should be a number

OpenStudy (anonymous):

lets start at this line here \[y'(2y+\frac{1}{y})=-\frac{1}{x}\] now \(x=e,y=1\) so you get \[3y'=-\frac{1}{e}\] and therefore \(y'=-\frac{1}{3e}\) is your slope, assuming my algebra was right i is a number, there is no variable in the slope

OpenStudy (anonymous):

Alright, I'll go with it, thank you so much for the help!